# On the Use of Large Language Models in Mutation Testing

### v2

@@ 3.3.2. Prompts
Figure 1. The Default Prompt Template
3.4. Evaluation Metrics
To assess the utility of the LLM-generated mutations, we should define a set of metrics that comprehensively evaluate them. Therefore, we designed several categories of metrics capturing different aspects as follows: Cost Metric, Usability Metrics, and Behavior Metrics. Besides these metrics, we also additionally report the Mutation Count and Mutation Score. Note that both of them cannot directly reflect the quality of generated mutations. For example, a higher mutation score does not indicate a higher quality, which may also be caused by generating a large proportion of mutants that are easily killed.
To better understand the metrics, we first illustrate the categories of mutations. Let all generated mutations be the set $A$, where LLMs may generate syntactically incorrect mutations that can not pass compilation. Thus we denote the compilable mutations as the set $C$. Note that the set $C$ contains two types of mutations that should be removed, the useless mutations (denoted as $U$) and equivalent mutations (denoted as $E$). The useless mutations refer to those that are identical to the original code or other mutations. In mutation testing, the ideal set of mutations to employ is from the set $(C-E-U)$, excluding equivalent mutations that have no contributions to discovering the weakness of tests. However, determining the equivalent mutations set $E$ is undecidable, so we follow the common practice and use the set $(C-U)$ as a practical approximation to perform mutation testing (Diniz et al., 2021; Tian et al., 2022; Zhang et al., 2022).
3.4.1. Cost Metric
The Cost Metric indicates time costs in mutation generation.
@@ 3.4.1. Cost Metric
Average Generation Time:  the average duration needed to produce each mutation, measuring the efficiency of an approach.
3.4.2. Usability Metrics
The Usability Metrics serve as key indicators of the practical utility of generated mutations.
Compilability Rate:  the proportion of mutations that successfully compile, computed as: ${|C|}/{|A|}$. Because LLMs cannot guarantee the generation of entirely correct code and the compilation process can be time-consuming, we prefer the approach with a higher comparability rate.
Useless Mutation Rate:  the proportion of useless mutations, computed as: $|U|/|A|$. LLMs sometimes generate duplicate mutations or mutations that are exactly the same as the original code. While traditional approaches rarely generate useless mutations, LLMs suffer from issues of misunderstanding and repetition, leading to generating these useless mutations.
Equivalent Mutation Rate:  the proportion of equivalent mutation, computed as: $|E|/|A|$. Equivalent mutations are semantically equivalent to the original program, which impacts the accuracy of the mutation score. Due to the undecidable nature of identifying whether a mutation is equivalent, we sample significant portions of all mutations and judge them manually. More specifically, given the population to be the set $A$, we set the confidence level to be 95% and the margin of error to be 5%, to get a subset of mutations to be manually inspected.
AST Node Diversity:  We prefer the set of mutations with diverse forms, therefore, we parse the code before and after mutation, and check the diversity of newly introduced AST nodes. The approach with a greater variety of AST node types is considered superior.
3.4.3. Behavior Similarity with Real Bugs
Behavior similarity refers to the degree of similarity in functionality and execution between a mutation and its corresponding real bug. Because capturing all program behaviors is undecidable, following existing studies, we thus employ the following metrics as a proxy to measure behavior similarity (Ojdanic et al., 2023a; Papadakis et al., 2018; Khanfir et al., 2023).
Real Bug Detectability:  Given the location of a real bug (which may occupy multiple lines), let a mutation generation approach generate a set of mutations, if the tests that kill these mutations can also kill this real bug, we say that the mutation generation approach can “detect” this real bug. Given a set of real bugs and a mutation generation approach, Real Bug Detectability quantifies the number of bugs detected by this mutation approach.
Coupling Rate:  Mutation testing is based on the fault coupling effect, i.e., mutations though are simple syntactic changes, coupled to more complex bugs (Offutt, 1992). To describe the extent to which a mutation couples to its corresponding real bug, we adopt the Coupling Rate following the existing study (Just et al., 2014b). Given a real bug and a corresponding mutant generated at the same location, the mutant is deemed a coupled mutant if the set of the test cases killing the mutant overlaps with the set of the bug-triggering test cases. The Coupling Rate refers to the proportion of coupled mutants within $(C-U)$.
Ochiai Coefficient:  approximates the semantic similarity by passing and failing test cases (Offutt and Untch, 2001), which is also widely used in the field of fault localization. Following the existing study (Ojdanic et al., 2023a), we employ the Ochiai Coefficient to calculate how a mutation resembles its corresponding real bug. More concretely, let $M$ and $B$ be a mutation and its corresponding real bug, and let $fT_{M}$ and $fT_{B}$ be the set of their failing test cases respectively. $M$’ coefficient is computed by the following formula:
(1) $Ochiai(M,B)=\frac{|fT_{M}\cap fT_{B}|}{\sqrt{|fT_{M}|\times|fT_{B}|}}$
Therefore, if $M$’s Ochiai Coefficient is $1$, it is semantically equal to its corresponding bug $B$.
3.5. Experiment Settings
3.5.1. Settings for Mutation Generation
One major target of our study is to investigate the similarity between LLM-generated mutations and real bugs. Therefore, we generate mutations within the context of real bugs to enable such a comparison. To determine the number of mutations generated in a prompt, we adopt the setting of PIT (Coles et al., 2016). We use PIT-generated mutations with the full set of operators and find that PIT generates 1.15 mutations per line of code. Therefore, based on the given code context, we generate one mutation per line. We explore the influence of context length on mutation generation in Section 5.
3.5.2. Settings for RQ1-RQ3
In the first part of our study, we intend to qualitatively and comparatively evaluate the capability of LLMs in generating mutations. We adopt the default settings, which contain all six LLMs, and use the default prompt template, generating mutations on the universal set of the 851 bugs in Section 3.2.
To understand the performance of the LLMs, we adopt LEAM (Tian et al., 2022), $\mu$Bert (Degiovanni and Papadakis, 2022), PIT (Coles et al., 2016), and Major (Just et al., 2011) as baselines. Among these approaches, LEAM is the state-of-the-art mutation generation approach which employs a model trained from 13 million real Java bugs. While $\mu$Bert is based on Bert (Devlin et al., 2018), a masked pre-trained language model that can not conversational respond to human-like text and code.PIT (Coles et al., 2016) and Major (Just et al., 2011) are the popular traditional mutation testing tools with human-predefined operators. Note that for PIT and Major, we employ all mutation operators.
3.5.3. Settings for RQ4
In the second part of our study, we intend to explore how different prompts and different models impact the performance of mutation generation. Due to budget limitations, we sample a subset of 306 bugs from our dataset, which consists of 10 sampled bugs from each Defects4J project and all ConDefects bugs.
In RQ4, for the prompts exploration, we use two representative LLMs, i.e., the closed-source GPT-3.5-Turbo and the open-source DeepSeek-Coder-V2-236b, then switch different prompt templates. More specifically, we modify the default prompt template by adding and removing different sources of information in the Context, and we finally get 4 types of prompts, shown as follows:

### v3

@@ 2.5.2. Mutation Generation via LLMs
3. Evaluation Results
In this section, we aim to comprehensively evaluate these approaches via various aspects. Table 4 presents the overall performance of each generation approach. In terms of the number of mutations generated (Mutation Count), PIT produces the largest number of mutations at 340,728, significantly more than other approaches, while the remaining approaches range from about twenty thousand to thirty thousand. In terms of mutation score, the traditional rule-based approaches (PIT and Major) are over 0.5, LEAM and $\mu$Bert are over 0.6, and the LLMs from the GPT family are all over 0.7.
Note that the number of generated mutations and mutation scores are not quality metrics for evaluating these approaches. A large number of mutations may include redundant or equivalent mutants that do not contribute to mutation scores, while a high mutation score may be influenced by the ease with which mutants are detected rather than their resemblance to real faults.
Metric PIT Major LEAM $\mu$BERT GPT-3.5 GPT-4o GPT-4o-M DS-236b SC-16b CL-13b
Mut. # 340728 20831 36505 12946 31656 31945 34417 34741 28058 27614
Mut. Score 0.541 0.521 0.648 0.677 0.702 0.737 0.708 0.714 0.507 0.669
Real Bug Detec. 40.1% 66.9% 63.9% 60.5% 87.0% 89.9% 90.8% 89.4% 38.5% 71.2%
Avg. Ochiai 23.0% 35.5% 32.4% 29.3% 47.7% 56.2% 54.1% 50.9% 26.0% 33.3%
Coupling Rate 23.5% 38.9% 38.1% 47.3% 46.9% 50.6% 49.5% 49.5% 34.0% 45.6%
Comp. Rate 100% 97.6% 44.5% 26.5% 62.5% 76.4% 75.3% 75.8% 16.4% 69.3%
Dup. Mut. Rate 0.0% 0.0% 2.3% 1.7% 10.3% 7.8% 7.0% 7.7% 12.8% 37.1%
A. G. Time (s) 0.02 0.08 3.06 2.34 1.79 1.31 1.65 4.25 7.53 9.06
Note: The mutation generation approach achieving the best performance in terms of a metric is highlighted in grey background color. GPT-4o-M: GPT-4o-Mini. DS-236b: DeepSeek-Coder-V2-236b. SC-16b: StarChat-$\beta$-16b. CL-13b: CodeLlama-Instruct-13b.
3.1. RQ1: Performance on Effectiveness
To measure the effectiveness of mutation generation approaches, we evaluate them with three sub-RQs, which estimate their effectiveness via real bug detectability, coupling, and diversity, respectively.
3.1.1. RQ1.1 Detectability
As illustrated in Section 2.4.1, we estimate real bug detectability via the following metric.
Real Bug Real Bug Detection Rate: The row Real Bug Detec. of Table 4 presents the proportion of real bugs that can be detected across different mutation generation approaches. Among the traditional approaches, PIT achieves a detection rate of 40.1%, while Major, LEAM, and $\mu$Bert perform significantly better with a detection rate of 66.9%, 63.9%, and 60.5%, respectively. For LLM-based approaches, the results are significantly more effective than those of traditional ones. GPT-4o-Mini achieves the highest detection rate at 90.8%, closely followed by GPT-4o and DeepSeek, with detection rates of 89.9% and 89.4%, respectively. Meanwhile, GPT-3.5 and CodeLlama achieve moderate detection rates of 87.0% and 71.2%, respectively.
Table 5. Number of Real Bugs Detected by Each Mutation Generation Approach
Project Bug # PIT Major LEAM $\mu$BERT GPT-3.5 GPT-4o GPT-4o-M DS-236b SC-16b CL-13b
Chart 26 19 17 16 10 22 22 23 23 9 8
Lang 65 23 51 27 38 52 55 56 57 11 41
Time 27 12 23 18 15 25 24 26 22 10 17
Math 106 40 82 44 65 92 94 96 95 58 77
Mockito 38 14 28 27 5 30 29 36 34 16 21
Closure 133 38 92 122 103 130 131 131 132 79 114
Cli 39 6 11 27 29 37 39 39 39 2 34
Codec 18 6 11 12 11 14 16 17 17 1 11
Csv 16 7 13 13 7 15 16 16 16 2 13
Gson 18 6 9 11 12 13 16 16 15 1 10
JacksonCore 26 1 17 17 18 23 23 21 21 1 17
Jsoup 93 48 77 90 86 93 92 93 90 57 87
ConDefects 246 121 138 120 116 194 208 203 200 81 156
ALL 851 341 569 544 515 740 765 773 761 328 606
Note: The mutation generation approaches that achieve the best performance in one project are highlighted in grey background color.
To better understand the detection of the generation approaches, we analyze their project-wise detectability, shown as Table 5. For individual projects, LLM-based approaches consistently achieve top performance across multiple projects. Particularly, GPT-4o-Mini outperforms other approaches in several projects, including Chart, Time, Mockito, and Jsoup, where it detects the highest number of bugs. Similarly, DeepSeek and GPT-4o demonstrate strong performance, achieving the best results in projects such as Lang, Math, and Closure.
3.1.2. RQ1.2 Coupling
To assess the degree of coupling, we perform comprehensive experiments that include both an overall metric evaluation and a project-wise analysis.
Average Ochiai Coefficient: The row Avg. Ochiai of Table 4 presents the Ochiai Coefficient values across the approaches, which measures the semantic distance between mutations and real bugs. Among the traditional approaches, Major achieves an Ochiai coefficient of 35.5%, higher than the others. LEAM and $\mu$BERT achieve coefficients of 32.4% and 29.3%, respectively, indicating that the best-performing rule-based approach, Major, outperforms both of them. The LLM-based approaches demonstrate significantly higher Ochiai coefficients, with GPT-4o achieving the highest value of 56.2%, followed closely by GPT-4o-Mini at 54.1%, and DeepSeek at 50.9%. Overall, GPT-4o and GPT-4o-Mini exhibit their superiority in generating mutations with smaller semantic distances with real bugs.
High Ochiai Bug Count: Following the study of Ojdanic et al. (Ojdanic et al., 2023), the mutation set with an Ochiai coefficient over 0.8 is considered to be of high semantic similarity with its corresponding real bug. For each project, we count the number of bugs that achieved a high Ochiai coefficient by each approach, shown as Table 6. The overall trend is similar with Real Bug Detectability, which indicates that the LLM-based appraoches, GPT-4o and GPT-4o-Mini, achieve the highest number of bugs across most projects. DeepSeek also exhibits strong performance, closely following GPT-4o in many cases. All traditional approaches exhibit significantly fewer high Ochiai bugs. Interestingly, among traditional approaches, Major outperforms LEAM and $\mu$BERT across most projects.
Table 6. Number of Bugs with High Ochiai Scores (Over 0.8) Achieved by Each Mutation Generation Approach
Project Bug # PIT Major LEAM $\mu$BERT GPT-3.5 GPT-4o GPT-4o-M DS-236b SC-16b CL-13b
Chart 26 9 9 3 5 10 13 12 12 7 3
Lang 65 15 34 13 18 32 35 37 34 12 23
Time 27 8 14 8 4 8 12 12 11 4 7
Math 106 20 42 13 27 46 56 48 47 19 33
Mockito 38 1 12 5 3 14 13 17 14 3 3
Closure 133 12 25 40 14 38 69 56 45 21 26
Cli 39 3 3 13 8 18 22 23 20 4 10
Codec 18 3 1 3 3 2 8 6 7 3 2
Csv 16 4 2 1 2 4 10 10 8 3 1
Gson 18 2 1 4 3 2 8 7 6 3 2
JacksonCore 26 2 3 9 8 11 14 12 11 5 6
Jsoup 93 16 6 32 20 46 39 39 41 25 19
ConDefects 246 101 150 132 134 175 179 181 177 112 148
ALL 851 196 302 276 249 406 478 460 433 221 283
Note: The mutation generation approaches that achieve the best performance in one project are highlighted in grey background color.
Coupling Rate: The overall Coupling Rate for each mutation generation approach is shown in the last row of Table 4. The traditional approaches, PIT, Major, LEAM, and $\mu$Bert, exhibit lower coupling rates than LLM-based approaches, with coupling rates of 23.5%, 38.9%, 38.1%, and 47.3%, respectively. Among the LLM-based approaches, GPT-4o achieves the highest coupling rate, with a value of 50.6%. It is followed closely by GPT-4o-Mini and DeepSeek, with coupling rates of 49.5% and 49.5%, respectively. Overall, the trends are similar to the behavior similarity discussed in Section 3.1, indicating the superiority of the LLM-based approaches, particularly GPT-4o, in generating highly coupled mutations that better mimic real bugs.
Table 7. Distribution of Coupling Categories Across Mutation Generation Approaches
Category PIT Major LEAM $\mu$BERT GPT-3.5 GPT-4o GPT-4o-M DS-236b SC-16b CL-13b
Strong 10.1% 9.6% 4.9% 7.8% 7.0% 6.8% 6.2% 6.4% 8.0% 6.8%
Strong Addi. 4.7% 5.8% 4.3% 5.2% 6.7% 10.2% 12.1% 10.3% 5.7% 6.0%
Partial 1.8% 1.1% 0.5% 0.8% 1.2% 1.2% 0.8% 1.0% 2.7% 0.9%
Partial Addi. 6.9% 22.4% 28.4% 33.5% 32.0% 32.5% 30.4% 31.8% 17.6% 31.9%
Not Subs. 30.6% 13.2% 26.7% 20.4% 23.3% 23.1% 21.3% 21.9% 16.7% 21.3%
Not Detected 45.9% 47.9% 35.2% 32.3% 29.8% 26.3% 29.2% 28.6% 49.3% 33.1%
Note: Strong - Strong Substitution, Strong Addi. - Strong Substitution + Additional Failing Tests, Partial - Partial Substitution, Partial Addi. - Partial Substitution + Additional Failing Tests, Not Subs. - Not Substitution.
Coupling Categorization: As mentioned in Section 2.4.1, we classify each mutation of each generation approach into different scales, i.e., Strong Substitution, Strong + Additional Tests Fail, Partial Substitution, Partial + Additional Tests Fail, Not Substitution, and Not Detected. Note that the first four categories are used to calculate the Coupling Rate. Table 7 presents the distribution of each generation approach. Among the traditional approaches, PIT and Major exhibit a significantly higher proportion of mutants in the Not Detected category, with rates of 45.9% and 47.9%, respectively, indicating that a substantial number of their generated mutants fail to couple with real bugs. However, they achieve relatively higher proportions in the Strong category, with PIT at 10.1% and Major at 9.6%, suggesting that while their coupling ability is limited, they can still generate a small number of strongly coupled mutations. In contrast, LLM-based approaches such as GPT-4o and GPT-4o-Mini, demonstrate stronger coupling capabilities across multiple categories. These approaches achieve significantly higher rates in the Partial Substitution + Additional Tests Fail category, with values of 32.5% and 30.4%, respectively. Furthermore, they outperform traditional methods in the Strong Substitution + Additional Tests Fail category, with GPT-4o and GPT-4o-Mini achieving 10.2% and 12.1%, respectively. Overall, while the traditional rule-based approaches generate a larger proportion of strong substitution mutations, LLM-based approaches demonstrate a stronger ability to generate diverse and highly coupled mutations.
3.1.3. RQ1.3 Diversity
We explore the syntactic diversity of mutations generated by each approach. Note that PIT works at the Java byte code level, we thus skip it in this RQ.
Figure 2. Number of Newly Introduced AST Nodes of the Studied Approaches and the Few-Shot Examples
AST Node Diversity: To assess the diversity of AST node types, we first determine if a mutation involves deletion (i.e., removing a code element). For the non-deletion mutations, we parse the code before and after mutation by the parser Javalang [^3] https://pypi.org/project/javalang, to observe which new AST node types are introduced. A larger number of newly introduced AST node types indicates richer diversity.
Figure 2 illustrates the number of introduced AST node types, with traditional approaches highlighted in green and LLM-based approaches in blue. The figure reveals that LLM-based mutation approaches introduce more types of new AST nodes than traditional approaches. The best LLM-based approach, DeepSeek, exhibits the greatest diversity, introducing 49 new AST node types, followed by GPT-4o-Mini, which introduces 45 new types. Additionally, all six LLMs are significantly higher than all other approaches in AST node types. We also check the AST node diversity of the examples we provided in the LLM prompts (see Table 3), as shown by the “Examples” in Figure 2. Interestingly, we find that these examples introduce only one new type of AST node and are subsumed by all other approaches. This indicates that LLMs are creative in generating mutations, producing significantly more diverse mutations than the examples provided to them.
Table 8. Ratio of Deletion Mutations and the Top-3 Newly Introduced AST Node Types for Each Approach
Top K Major LEAM $\mu$BERT GPT-3.5 GPT-4o GPT-4o-M DS-236b SC-16b CL-13b
Del. 15.3% 48.9% 0.4% 5.3% 2.1% 0.5% 1.0% 0.3% 0.1%
1st ST(78.6%) LT(23.2%) LT(29.4%) LT(39.2%) LT(45.4%) LT(43.3%) LT(38.8%) LT(22.7%) LT(47.8%)
2nd LT(21.4%) MR(14.4%) SE(25.7%) BO(17.3%) BO(16.6%) BO(21.6%) BO(19.1%) BO(13.8%) BO(25.7%)
3rd —— CT(9.6%) MR(24.5%) MI(9.9%) CC(8.8%) CC(7.6%) MI(10.1%) MR(13.1%) MI(9.3%)
Del.: The proportion of deletion mutations. LT: Literal, BO: BinaryOperation, MI: MethodInvocation, MR: MemberReference, SE: StatementExpression, ST: Statement.
We further analyzed the distribution of deletions and AST node types across different approaches, as shown in Table 8, which highlights the top three most frequent AST node types. For the proportion of deletions, the traditional approach, LEAM, shows the highest deletion mutation ratio, with 48.9%, followed by Major at 15.3%. In contrast, all LLM-based approaches generate deletion mutations at a rate below 6%, with three models producing less than 1.0%. This indicates that, compared to existing approaches, LLMs are less inclined to generate deletion mutations. As shown later, we find that deletions are one of the key factors contributing to the lower compilation success rate.
Regarding AST node types, the most frequent new AST node type of all the LLM-based approaches is Literal (e.g., $a>b\mapsto\text{true}$). However, for the traditional approach Major, the most frequent node type is Statement (e.g., replacing a statement with an empty statement). Note that although these mutations are not with the deletion operator, a significant portion of the production of empty statements exists, leading to a large proportion.
AST Edit Distance: To quantify the extent of syntactic modifications introduced by each mutation generation approach, we compute the AST edit distance for each mutation using Gumtree (Falleri et al., 2014). The distribution of edit distances for each approach is presented in Table 9, providing insights into the complexity and diversity of the generated mutations.
Table 9. The Distribution of AST Edit Distance for Each Approach
Distance Major LEAM $\mu$BERT GPT-3.5 GPT-4o GPT-4o-M DS-236b SC-16b CL-13b
1 69.2% 58.5% 73.8% 52.4% 50.0% 47.3% 45.1% 36.0% 46.4%
2 27.8% 23.6% 8.9% 36.4% 38.8% 41.9% 44.2% 39.1% 46.6%
3 1.8% 7.3% 2.7% 6.8% 7.2% 6.3% 6.5% 8.0% 3.8%
4 0.6% 3.4% 14.2% 2.4% 2.2% 2.5% 2.6% 6.0% 1.8%
5 0.4% 1.0% 0.1% 0.9% 0.8% 0.8% 0.8% 3.7% 0.5%
$\geq$6 0.2% 6.1% 0.4% 1.0% 1.0% 1.2% 0.8% 7.2% 1.0%
From the table, we can find that all traditional approaches predominantly generate small-scale changes. Specifically, the proportion of mutations with an edit distance of 1 for Major, LEAM, and $\mu$Bert is 69.2%, 58.5%, and 73.8%, respectively, indicating they involve minimal syntactic changes to the original code in the majority of cases. In contrast, LLM-based approaches exhibit a broader range of edit distances, reflecting their capability to introduce more complex transformations. For example, GPT-4o exhibits a more balanced distribution, with 38.8% of its mutations having an edit distance of 2, 7.2% with a distance of 3, and 2.2% with a distance of 4, indicating it generates both minor and moderate code changes.
Answer to RQ1: LLM-based approaches, such as GPT-4o and GPT-4o-Mini, outperform traditional methods in effectiveness. The weighted averages of the rule-based approach in terms of real bug detection rate and coupling rate are 41.64% and 24.38%, while the LLMs are 79.07% ($\uparrow 89.89\%$; 37.43% PPs) and 46.21% ($\uparrow 89.54\%$; 21.83% PPs). Specifically, GPT-4o-Mini achieves the highest real bug detection rate (90.8%) compared to PIT (40.1%) and Major (66.9%). GPT-4o also achieves the highest coupling rate (49.9% vs. 17.9% for PIT and 42.7% for Major). Moreover, LLM-generated mutations demonstrate significantly higher syntactic diversity, with significantly higher new AST node types (e.g., DeepSeek has 49), and a lower proportion of deletions (e.g., 26.3% for GPT-4o vs. 47.9% for Major).
3.2. RQ2: Performance on Validity
To evaluate the validity of mutation generation approaches, we consider the aspects that may impact their validity, i.e., comparability rate, duplicate mutation rate, and equivalent mutation rate. This RQ is divided into the following sub-RQs.
3.2.1. RQ2.1: Compilability
The row Comp. Rate of Table 4 presents the overall Compilability Rate (CR) of each approach. Among the traditional approaches, the rule-based approaches, PIT and Major, demonstrate superior compilability, with PIT achieving a perfect 100% $CR$, and Major following closely at 97.60%. Note that although Major operates under simple syntax rules, it can still generate non-compilable mutants. For example, the Java compiler’s analyzer rejects constructs like while(true). PIT uses the mutation operators more restrictedly and previously filters out non-compilable mutations. In contrast, LEAM and $\mu$Bert, have rates at 35.0% and 22.5%, respectively. For LLM-based approaches, there is a considerable variation in compilability. GPT-4o achieves the highest compilability among LLMs at 76.40%, followed closely by GPT-4o-Mini and DeepSeek, with rates of 75.30% and 75.80%, respectively. Other LLMs such as GPT-3.5 and CodeLlama show moderate $CR$s of 62.50% and 71.20%, respectively, whereas models like StarChat fall behind with a $CR$ of 38.50%. Compared to small models (i.e., LEAM and Bert), GPTs and DeepSeek show significantly higher compilation success rates. For example, the compilability rate of GPT-4o is over 30% higher than that of LEAM and $\mu$Bert.
3.2.2. RQ2.2: Duplicate Mutation Rate
As aforementioned, duplicate mutations refer to those that are redundant or identical to the original code, providing no meaningful contribution to the mutation testing process. Therefore, a lower Duplicate Mutation Rate ($DMR$) is desirable, as it indicates a more efficient mutation generation process with minimal redundancy. The row Dup. Mut. Rate of Table 4 presents the results.
Rule-based approaches, PIT and Major, achieve a perfect 0% $DMR$. This is expected, as these approaches apply well-defined mutation rules that ensure each generated mutant introduces a meaningful change to the program. Other traditional approaches, i.e., LEAM and $\mu$Bert, show relatively low $DMR$s of 2.30% and 1.70%, respectively. In contrast, LLMs generate a significant portion of duplicate mutations, exhibiting a higher $DMR$. For the LLMs from OpenAI, GPT-3.5, GPT-4o, and GPT-4o-Mini exhibit moderate $DMR$s of 10.30%, 7.80%, and 7.00%, respectively. The open-source LLMs such as StarCoder and DeepSeek demonstrate higher $DMR$s of 12.80% and 7.70%, respectively, while CodeLlama shows the highest $DMR$ at 37.10%.
3.2.3. RQ2.3: Equivalent Mutations
Given the undecidable nature of identifying equivalent mutations, we manually evaluate each mutation’s equivalence, following the methodology of the existing study (Kushigian et al., 2024). For each mutation generation approach, we randomly sample a subset of compilable mutations, ensuring a 95% confidence level with a 10% margin of error.
Two authors independently evaluate the equivalence of the selected mutations to minimize subjectivity. The procedure for checking a mutation involves first transplanting it into the original project and executing all tests. If it passes all tests, the author then manually verifies whether the mutation is semantically equivalent. Note that PIT does not generate source-level mutations, so its mutations must be manually extracted from Java bytecode and transformed back into source-level representations, which is significantly more time-consuming.
In total, 959 mutations were labeled, 508 of which were filtered by tests (i.e., killed), and the remaining 451 were manually checked. The entire process took approximately ten hours per person. The manual verification requires checking the code context, which may involve multiple classes and methods. For example, to check a mutation that replaces an integer with a method invocation, where the method is invoked by the this object but defined in the parent class, we should review the parent class file and verify whether its return value is consistently equal to the integer. We measure the consistency between the two authors by Cohen’s kappa coefficient ($\kappa$(Viera et al., 2005), achieving an overall $\kappa$ value of 0.8, which indicates a high level of agreement. In cases of disagreement, both authors conducted thorough discussions to resolve contradictions and finally reach consistent labels. The labeling results are available at our repo.
Table 10. Sample Compilable Mutations for Identifying Equivalent Mutations
Metric PIT Major LEAM $\mu$Bert GPT-3.5 GPT-4o GPT-4o-Mini DS-236b SC-16b CL-13b
#Sampled 97 98 96 94 96 96 96 96 94 96
#Test Filtered 26 52 48 53 60 64 59 63 32 51
#Eq. Mut 1 2 2 6 4 3 3 5 10 6
EMR 1.0% 2.0% 2.1% 6.4% 4.2% 3.1% 3.1% 5.2% 10.6% 6.3%
Table 10 presents the estimated Equivalent Mutation Rate ($EMR$), reporting the number of sampled mutations (#Sampled), the number of equivalent mutations we identified (#Eq. Mut), and the $EMR$ for each approach. Among the traditional approaches, PIT achieves the lowest $EMR$ of 1.0%, followed by Major with an $EMR$ of 2.0%, indicating that rule-based techniques produce fewer equivalent mutations. In contrast, the LLM-based approaches exhibit higher $EMR$s, with variations across different models. StarChat has the highest $EMR$s of 10.6%, while GPT-4o, GPT-4o-Mini, and DeepSeek show relatively lower $EMR$s, ranging from 3.1% to 5.2%. Overall, although the OpenAI models such as GPT-4o and GPT-4o-Mini, tend to have slightly higher $EMR$s than PIT and Major, the increase remains modest and within an acceptable range.
Note that the estimated $EMR$ of Major given by the recent study of Kushigian et al. (Kushigian et al., 2024) is 2.97%, where the authors sampled a subset of mutation for manual judgment. We adopt the same experimental process and the result of Major is 2.0%, which is consistent with theirs considering the randomness introduced by sampling.
Answer to RQ2: Rule-based approaches (i.e., PIT and Major) outperform LLMs in comparability rate, duplicate mutation rate, and equivalent mutation rate, with a weighted average of 99.90%, 0.00%, and 1.10% respectively. While the weighted average of LLMs is 63.80% ($\downarrow 36.14\%$; $36.10\%$ PPs), 13.10% ($\uparrow 13.10\%$; $13.10\%$), and 5.30% ($\uparrow 79.25\%$; $4.2\%$ PPs), respectively. Specifically, PIT and Major produce few non-compilable or duplicate mutations, while GPT-4o generates 23.6% non-compilable mutations and 7.8% duplicate mutations. Although rule-based approaches are slightly better in terms of equivalent mutation rates, the equivalent mutation rates across approaches remain comparable, with no significant disparities observed.
3.3. RQ3: Performance on Efficiency
@@
To evaluate the performance on efficiency, we record the average generation time for each approach.
Average Generation Time: The row A. G. Time of Table 4 presents the results of Average Generation Time. Among the traditional appraoches, PIT demonstrates the fastest generation time at 0.02 seconds per mutation, followed by Major, which takes 0.08 seconds, making them the most efficient approaches. In contrast, LEAM and $\mu$Bert take 3.06 seconds and 2.34 seconds, respectively. Among the LLMs, the LLMs of OpenAI (i.e., GPT-3.5, GPT-4o, and GPT-4o-Mini) show relatively faster generation, ranging from 1.31 to 1.79 seconds. While the open-source LLMs StarChat, CodeLlama, and DeepSeek require 7.53, 9.06, and 4.25 seconds, respectively. The results suggest that the rule-based approaches are the most time-efficient for generating mutations, which are tens to hundreds of times faster than other approaches.
Answer to RQ3: Rule-based traditional approaches, such as PIT and Major, outperform LLMs in time cost. On average, PIT and Major generate a single mutant in just 0.02s and 0.08s, respectively, whereas the best-performing LLM, GPT-4o, requires 1.31s.
3.4. RQ4: Categorizations of Non-Compilable Mutations
To answer this RQ, we first identify the error types of non-compilable mutations and then analyze the surrounding code context types that are error-prone to generate them.
3.4.1. Categorizations of compilation error types
Non-compilable mutations require a compilation step to filter out, which results in wasted computational resources. As mentioned in Section 3.2, LLMs generate many non-compilable mutations. This RQ analyzes the types of errors and potential root causes of these non-compilable mutations. Following the setting of the previous steps, we first sample 384 non-compilation mutations from the outputs of GPT models, ensuring the confidence level is 95% and the margin of error is 5% (Linares-Vásquez et al., 2017; Tufano et al., 2019; Macklon et al., 2023; de Mello et al., 2015). From the manual analysis of these non-compilable mutations, we identified nine distinct error types, shown as follows.